# Optimizing Integrated Terrestrial and Non-Terrestrial Networks Performance with Traffic-Aware Resource Management

### v1

@@ Abstract
To address an ever-increasing demand for ubiquitous high-speed connectivity, mobile networks have intensified their deployment process. However, achieving this target has proven to be a challenge and has led to a surge in overall energy consumption. In recent years, non-terrestrial networks have been endorsed as a potential solution to these problems by complementing the coverage of the terrestrial network in areas with limited network deployment. To this end, this paper proposes an integrated terrestrial and non-terrestrial network (TN-NTN) that utilises the overall available communication resources to expand coverage and meet Quality of Service (QoS) requirements during high-traffic hours in any deployment scenario. Importantly, our framework allows to drastically reduce the terrestrial network energy consumption during low-traffic hours. Specifically, we introduce a novel radio resource management algorithm, BLASTER (Bandwidth SpLit, User ASsociation, and PowEr ContRol), which integrates bandwidth allocation, user equipment (UE) association, power control, and base station activation within the TN-NTN. This algorithm aims to optimize network resource allocation fairness and energy consumption dynamically, demonstrating new opportunities in deploying satellite networks in legacy cellular systems. Our study offers a comprehensive analysis of the integrated network model, emphasizing the effective balance between energy saving and QoS, and proposing practical solutions to meet the fluctuating traffic demands of cellular networks.
@@ I Introduction
In recent years, the rapid advancement of cellular communications has driven a marked increase in the demand for high-speed data connectivity. This surge has led to stringent requirements for achieving widespread network connectivity and delivering enhanced network capacity. To meet these challenges, mobile operators have intensified terrestrial macro base stations deployment. However, this approach has its limits, as coverage cannot be guaranteed in logistically challenging locations [1]. In addition, the extensive deployment of terrestrial networks with ever-increasing capabilities to deliver high capacity also escalates the overall network energy consumption—a concern in light of current environmental and economic conditions. Consequently, a principal aim in the development and management of mobile networks is minimizing energy consumption, while adhering to QoS standards [2]. Fortunately, recent advancements related to NTNs have offered a promising alternative for extending coverage and enhancing capacity. NTNs utilise airborne vehicles such as unmanned aerial vehicles, high-altitude platform stations (HAPS), or satellites, which serve as MBSs or relay nodes to enable connectivity for UEs across the network. The primary benefit of NTNs is their capability to provide broad coverage over large and/or remote areas, where establishing terrestrial MBSs would be either too expensive or difficult. Among the different options for deployment, low-earth orbit (LEO) satellites are poised to become the leading method to achieve high-capacity connectivity from space [3, 4]. In fact, orbiting at altitudes between $200$ and $2000$ kilometers, their relative proximity to Earth provides enhanced signal strength and lower latency compared to other types of constellations. Consequently, this proximity leads to reduced energy needs for launching and lower power usage for transmitting signals to and from the satellite [4]. Spearheaded by companies like SpaceX and OneWeb, ongoing LEO constellation projects aim to launch thousands of LEO satellites around the Earth to establish an ultra-dense constellation. Through a collaboration with mobile operators, the objective is to create an integrated TN-NTN, which can deliver seamless and high-capacity communication services [5], as well as guarantee efficient services for UEs in the future [6, 7].
@@ I-A Related works
The resource allocation problem in an integrated TN-NTN has garnered some interest in the last few years. The authors in [8] have introduced a novel radio resource management scheme designed to enhance the performance of an integrated TN-NTN system using 5G 3rd Generation Partnership Project (3GPP) new radio (NR). The proposed solution leverages multicast subgrouping techniques to group UEs within the coverage areas of both terrestrial and non-terrestrial MBSs, ensuring optimal resource allocation to boost overall network throughput. The aim was to reduce the UE ping-pong effect between both tiers, increase resilience to frequent handovers, and simultaneously improve the quality of service perceived by UEs located at cell edges, ultimately enhancing their data rate experience.
@@
Load Balancing in terrestrial networks is a topic that has been well studied over the past few years, but in the integrated TN-NTN scenario, the contributions are limited. Typical methods used for load balancing involve the optimization of a utility function through a pricing-based association strategy [11, 12]. In this line, the authors in [7] have examined an integrated TN-NTN set up in an urban setting and have shown that diverting some of the traffic to LEO satellites improves the overall signal quality and decreases outages accordingly.
In [13] and [14], the authors have investigated the uplink performance of an integrated TN-NTN, leveraging LEO satellites to provide backhaul support to terrestrial MBSs. Both studies aimed to maximise the total uplink data rate, while adhering to backhaul capacity limits. To do this, [14] has taken into account minimum rate requirements and has adjusted the bandwidth division between fronthaul and backhaul links, while [13] has enhanced UE association and power management using matching algorithms.
@@
As stated previously, in the current environmental context, improving network energy efficiency (EE) and reducing energy consumption have become major objectives. To this end, mobile operators need to adapt the offered capacity by the terrestrial MBSs to the rate requirements by dynamically adjusting the number of active MBSs. For example, operating all MBSs during low-traffic periods is suboptimal, as keeping underused or idle terrestrial MBSs active leads to wasteful use of energy and communication resources [16]. In the context of an integrated TN-NTN, it is also beneficial to deactivate some terrestrial MBSs and handover the UEs to satellites to decrease the energy consumption of the terrestrial network.
MBSs activation is a well-studied topic in terrestrial networks. The authors of [17] have introduced an energy-efficient algorithm that strategically shuts down MBSs one at a time, ensuring they do not overburden neighbouring MBSs. To preserve the QoS, [18] has examined the effects of traffic offloading in heterogeneous networks on energy use and proposed a centralized Q-learning method to strike a balance between energy saving and QoS satisfaction. In addition, the authors of [19] have devised an algorithm that enables UEs to associate with multiple MBSs across different frequency bands, simultaneously optimising the transmit power of the MBSs to facilitate their shutdown during periods of low traffic. The authors of [20] have also tackled traffic uncertainties in ultra-dense networks by optimising both MBS activation and UE association strategies, employing chance constraint programming based on statistical traffic data to effectively balance traffic loads and reduce interference.
To the best of our knowledge, however, only the authors of [21] and [22] have considered NTNs as a solution for meeting coverage and capacity demands while deactivating MBSs. In more detail, [21] has investigated the challenges of MBSs activation in an integrated TN-NTN using HAPS. The authors have focused on offloading traffic from deactivated terrestrial MBSs to the HAPS, mainly using a sorting algorithm, which prioritises switching off MBSs with relatively lower traffic loads. In our latest work [22], we have introduced a framework which leverages the large coverage ability of LEO satellites to shutdown terrestrial MBSs in rural areas during low traffic.
@@ I-B Contributions
In this paper, we consider an integrated TN-NTN and introduce a novel radio resource management algorithm that adapts to the fluctuating traffic of the network.
The contributions of this paper are summarized as follows:
  • We develop BLASTER (Bandwidth SpLit, User ASsociaTion, and PowEr ContRol), an innovative approach to adaptive radio resource management. BLASTER seamlessly integrates the management between terrestrial and non-terrestrial tiers of bandwidth allocation and UE association. It also controls the transmission power and activation of terrestrial MBSs. The proposed framework is designed to balance network fairness and energy consumption by adjusting to current traffic conditions. Our findings demonstrate that BLASTER can reduce the overall terrestrial network energy usage by $67\,\%$ when compared to a standard integrated TN-NTN system adhering to 3GPP guidelines, while notably enhancing the average network sum log-throughput (SLT) in times of high traffic demand by at least $6\,\%$ compared to the same 3GPP benchmark mentioned above.
  • We propose a framework where the terrestrial and non-terrestrial tiers share the same spectrum. In this scenario, we demonstrate that the optimal bandwidth allocation for the non-terrestrial tier is proportional to the fraction of UEs associated to the satellites, provided that UEs have the same requirements. Through dynamic sharing of the resources for both tiers, we are able to optimize the collaboration of both tiers and enhance the network efficiency throughout the day.
  • By exploiting the special properties of the formulated optimization problem, we design a practical heuristic with an intuitive behaviour to solve the problem with a limited complexity while achieving results comparable to BLASTER.
@@ II System Model
Our study focuses on the downlink (DL) of a cellular network that consists of $M$ terrestrial MBSs and $N$ MBSs installed on a constellation of LEO satellites, making a total of $L$ MBSs. They provide service to $K$ UEs located in the area of study, which is composed of a rural and urban zone. We refer to the overall network bandwidth as $W$, which is distributed by the mobile network operator across the ground and space-based network. We suppose that the network operates within the S band, around 2 GHz, with ground and satellite MBSs using separate portions of this band. Throughout this paper, we will use $\mathcal{T}$ and $\mathcal{S}$ to represent the set of terrestrial and satellite MBSs respectively. Moreover, $\mathcal{B}=\mathcal{T}\cup\mathcal{S}=\{1,\cdots,j,\cdots,L\}$ is the complete set of MBSs, and $\mathcal{U}=\{1,\cdots,i,\cdots,K\}$ defines the set of UEs. For the remainder of the paper, we will designate the Hadamard product with $\odot$.
@@ II-A2 Satellite link channel model
It should be noted that components $SF_{ij}$ and $CL$ are function of -and greatly vary with- the LoS conditions, and that the LoS probability for the satellite links are provided in [24, Table 6.6.1-1]. Additionally, it should be highlighted that the elevation angle of the satellite also impacts the quality of the channel link. The elevation angle is the angle between the horizontal plane (the plane parallel to the surface of the Earth at the UE location) and the line of sight to the satellite. If the elevation angle increases, $CL$, $PL_{ij}^{e}$ and $SF_{ij}$ are greatly reduced. Given the cartesian coordinates of a satellite $s$ $\left(x_{s},y_{s},z_{s}\right)$ and a UE $u$ $\left(x_{u},y_{u},z_{u}\right)$, we can compute the elevation angle $\theta$ as:
$\theta=\arcsin\left(\frac{z_{s}-z_{u}}{\sqrt{\left(x_{s}-x_{u}\right)^{2}+\left(y_{s}-y_{u}\right)^{2}+\left(z_{s}-z_{u}\right)^{2}}}\right).$ (3)
@@
In Eq. (4), $p_{j}$ denotes the transmit power allocated per resource element (RE) at MBS $j$, $\mathcal{I}_{j}$ indicates the set of MBSs interfering with the serving MBS $j$ and $\sigma^{2}$ represents the noise power per RE. Moreover, assuming that MBS $j$ equally shares its total available bandwidth $W_{j}$ between the $k_{j}$ UEs it is serving, we are able to compute the mean throughput for UE $i$ connected to MBS $j$ as follows:
@@
II-C Power Consumption Model
The power consumption of a MBS can be modelled as the sum of multiple components [25]. Indeed, we denote as the baseline power consumption, the power used by the components that are kept active in a shut down MBS. Then, we denote as the static component, the power consumption that occurs regardless of the level of the MBS traffic load. The static power consumption represents the minimum power required to keep essential systems operational and maintain standby readiness. Finally, the dynamic component refers to the load-dependent power consumption that fluctuates depending on the MBS traffic load. Typically, the dynamic component increases whenever a MBS increases its transmit power or uses additional transmission resources, e.g., more resource blocks. For a MBS $j$, this model can be formulated as:
@@ II-C Power Consumption Model
where $P_{0}$ represents the baseline power consumption, $\psi_{j}$ represents the static component and $p_{j}$ accounts for the dynamic consumption of the MBS. Also, $||\cdot||_{0}$ is a binary-valued function equal to $1$ if the transmit power $p_{j}$ is greater than 0.
Importantly, note that we suppose the satellites in the constellation are solar-powered and their impact on their network power consumption can be neglected.

### v2

@@ Abstract
To address an ever-increasing demand for ubiquitous high-speed connectivity, mobile network deployments are becoming increasingly dense. However, this densification has also led to a surge in overall energy consumption, making the process increasingly challenging. In recent years, non-terrestrial networks have been mainly endorsed as a potential solution to enhance coverage by complementing the coverage of the terrestrial network (TN) in areas with limited network deployment. However, their ability to reduce TN energy consumption, though often overlooked, remains a significant advantage. To this end, this paper introduces a novel radio resource management algorithm, BLASTER (Bandwidth SpLit, User ASsociation, and PowEr ContRol), which integrates bandwidth allocation, user equipment (UE) association, power control, and base station activation within an integrated terrestrial and non-terrestrial network (TN-NTN). This algorithm aims to optimize network resource allocation fairness and energy consumption dynamically, demonstrating new opportunities in deploying satellite networks in legacy cellular systems. Our study offers a comprehensive analysis of the integrated network model, emphasizing the effective balance between energy saving and Quality of Service (QoS), and proposing practical solutions to meet the fluctuating traffic demands of cellular networks.
@@ I Introduction
In recent years, the rapid advancement of cellular communications has driven a marked increase in the demand for high-speed data connectivity. This surge has led to stringent requirements for achieving widespread network connectivity and delivering enhanced network capacity. To meet these challenges, mobile operators have intensified terrestrial macro base stations deployment. However, this approach has its limits, as coverage cannot be guaranteed in logistically challenging locations [1]. In addition, the extensive deployment of TNs with ever-increasing capabilities to deliver high capacity also escalates the overall network energy consumption—a concern in light of current environmental and economic conditions. Consequently, a principal aim in the development and management of mobile networks is minimizing energy consumption, while adhering to QoS standards [2]. Fortunately, recent advancements related to NTNs have offered a promising alternative for extending coverage and enhancing capacity. NTNs utilise airborne vehicles such as unmanned aerial vehicles, high-altitude platform station (HAPS), or satellites, which serve as MBSs or relay nodes to enable connectivity for UEs across the network. The primary benefit of NTNs is their capability to provide broad coverage over large and/or remote areas, where establishing terrestrial MBSs would be either too expensive or difficult. Among the different options for deployment, low-earth orbit (LEO) satellites are poised to become the leading method to achieve high-capacity connectivity from space [3, 4]. In fact, orbiting at altitudes between $200$ and $2000$ kilometers, their relative proximity to Earth provides enhanced signal strength and lower latency compared to other types of constellations. Consequently, this proximity leads to reduced energy needs for launching and lower power usage for transmitting signals to and from the satellite [4]. Spearheaded by companies like SpaceX and OneWeb, ongoing LEO constellation projects aim to launch thousands of LEO satellites around the Earth to establish an ultra-dense constellation. Through a collaboration with mobile operators, the objective is to create an integrated TN-NTN, which can deliver seamless and high-capacity communication services [5], as well as guarantee efficient services for UEs in the future [6, 7].
@@ I-A Related works
The resource allocation problem in an integrated TN-NTN has garnered some interest in the last few years. The authors in [8] have introduced a novel radio resource management scheme designed to enhance the performance of an integrated TN-NTN system using 5G 3rd Generation Partnership Project (3GPP) new radio (NR). The proposed solution leverages multicast subgrouping techniques to group UEs within the coverage areas of both terrestrial and non-terrestrial MBSs, ensuring optimal resource allocation to boost overall network performance. The aim was to reduce the UE ping-pong effect between both tiers, increase the resilience to frequent handovers, and simultaneously improve the quality of service perceived by UEs located at cell edges, ultimately enhancing their data rate experience.
@@
Indeed, interference management is a critical challenge to achieve efficient resource allocation, which must be addressed to ensure both tiers of the integrated TN-NTN can coexist and achieve optimal performance. To that end, studies such as [11, 12, 13] have provided important perspectives. The authors of [11] have developed a stochastic model to assess interference from terrestrial MBSs to satellite services in the upper $6$ GHz spectrum. The authors have introduced both stochastic and geometrical-stochastic interference models to evaluate the aggregated interference, ensuring it remains within acceptable limits for geostationary satellites. In [12], the authors have investigated the interference between terrestrial and non-terrestrial tiers in the $12$ GHz band, focusing on the potential impact of 5G MBSs on satellite systems. The authors have developed a simulation framework incorporating real-world deployment data and proposed strategies like exclusion zones and beamforming to mitigate interference and enable coexistence. Finally, [13] has presented a framework for spectrum sharing between beyond 5G TNs and satellite systems, optimizing MBS parameters such as transmit power and beam selection based on contextual factors like weather and satellite trajectory. The framework enhances spectrum utilization, while maintaining interference control, outperforming traditional, static spectrum-sharing policies.
Load Balancing in TNs is a topic that has also been well studied over the past few years, but in the integrated TN-NTN scenario, the contributions are limited. Typical methods used for load balancing involve the optimization of a utility function through a pricing-based association strategy [14, 15]. In this line, the authors in [7] have examined an integrated TN-NTN set up in an urban setting, and have shown that diverting some of the traffic to LEO satellites improves the overall signal quality and decreases outages accordingly. The authors of [16] and [17] have exploited the qualities of HAPSs to improve the QoS of ground UEs. [16] has proposed a fairness optimization approach for integrated TN-NTNs, using multiple-input multiple-output (MIMO) beamforming to improve spectral efficiency and manage interference, demonstrating superior performance over standalone TNs. In contrast, [17] has explored multi-connectivity offloading strategies using UAVs and HAPS in NTNs to reduce task computation latency for time-sensitive applications.
In [18] and [19], the authors have investigated the uplink performance of an integrated TN-NTN, leveraging LEO satellites to provide backhaul support to terrestrial MBSs. Both studies aimed to maximise the total uplink data rate, while adhering to backhaul capacity limits. To do this, [19] has taken into account minimum rate requirements, and has adjusted the bandwidth division between fronthaul and backhaul links, while [18] has enhanced UE association and power management using matching algorithms.
@@
As stated previously, in the current environmental context, improving network energy efficiency (EE) and reducing energy consumption have become major objectives. To this end, mobile operators need to adapt the offered capacity by the terrestrial MBSs to the rate requirements by dynamically adjusting the number of active MBSs. For example, operating all MBSs during low-traffic periods is suboptimal, as keeping underused or idle terrestrial MBSs active leads to wasteful use of energy and communication resources [2]. In the context of an integrated TN-NTN, it is also beneficial to deactivate some terrestrial MBSs and handover the UEs to satellites to decrease the energy consumption of the TN.
MBSs activation is a well-studied topic in TNs. The authors of [21] have introduced an energy-efficient algorithm that strategically shuts down MBSs one at a time, ensuring they do not overburden neighbouring MBSs. To preserve the QoS, [22] has examined the effects of traffic offloading in heterogeneous networks on energy use, and proposed a centralized Q-learning method to strike a balance between energy saving and QoS satisfaction. In addition, the authors of [23] have devised an algorithm that enables UEs to associate with multiple MBSs across different frequency bands, simultaneously optimising the transmit power of the MBSs to facilitate their shutdown during periods of low traffic. The authors of [24] have also tackled traffic uncertainties in ultra-dense networks by optimising both MBS activation and UE association strategies, employing chance constraint programming based on statistical traffic data to effectively balance traffic loads and reduce interference. Recent studies such as [25, 26, 27] have explored the integration of HAPS to enhance network efficiency in a more dynamic and sustainable manner. Indeed, [25] has investigated how HAPSs can complement traditional network densification to manage dynamic traffic in urban areas, demonstrating better energy efficiency and sustainability by using HAPSs to handle peak demand periods. In [26], the authors have tackled the traffic load estimation issue in HAPS-assisted networks, proposing Q-learning algorithms to optimize cell-switching strategies, improving energy efficiency and making advanced cell-switching methods feasible for vertical heterogeneous networks. The authors of [27] have investigated the challenges of MBSs activation in an integrated TN-NTN using HAPS. The authors have focused on offloading traffic from deactivated terrestrial MBSs to the HAPS, mainly using a sorting algorithm, which prioritises switching off MBSs with relatively lower traffic loads. Although the studies in [25, 26, 27] have provided valuable insights and promising results, they have not explored the optimization of spectrum sharing and allocation strategies, thereby limiting resource utilization efficiency in integrated TN-NTNs. In our latest work [28], we have introduced a framework which leverages the large coverage ability of LEO satellites to shutdown terrestrial MBSs in rural areas during low traffic. To the best of our knowledge, only the authors of [26], [27], and [28] have considered NTNs as a solution for meeting coverage and capacity demands while deactivating MBSs.
@@ I-B Contributions
In this paper, we consider an integrated TN-NTN, and introduce a novel radio resource management algorithm that adapts to the fluctuating traffic of the network. The contributions of this paper are summarized as follows:
  • We develop BLASTER (Bandwidth SpLit, User ASsociaTion, and PowEr ContRol), an innovative approach to adaptive radio resource management. BLASTER seamlessly integrates the management of bandwidth allocation and UE association between terrestrial and non-terrestrial tiers. It also controls the transmission power and activation of terrestrial MBSs. The proposed framework is designed to balance network fairness and energy consumption by adjusting to current traffic conditions. Our findings demonstrate that BLASTER can reduce the overall TN energy usage by $67\,\%$ when compared to a standard integrated TN-NTN system adhering to 3GPP guidelines, while notably enhancing the average network sum log-throughput (SLT) in times of high traffic demand by at least $6\,\%$ compared to the same 3GPP benchmark mentioned above.
  • We propose a framework where the terrestrial and non-terrestrial tiers orthogonally share the same frequency band based on the fluctuating traffic. In this scenario, we demonstrate that the optimal bandwidth allocation for the non-terrestrial tier is proportional to the fraction of UEs associated to the satellites, provided that UEs have the same requirements. Through dynamic allocation of the resources for both tiers, we are able to optimize the collaboration of both tiers and enhance the network efficiency throughout the day.
  • By exploiting the special properties of the formulated optimization problem, we design a practical heuristic with an intuitive behaviour to solve the problem with limited complexity, achieving results that highlight the trade-off between enhancing network SLT and reducing energy consumption.
Note that, since recent techno-economic [29, 30] studies have underscored the cost-effectiveness of the LEO satellites compared to HAPS, which are still suffering from technological, regulatory and economic constraints, we have focused on LEO satellites in this paper, but BLASTER is also applicable to HAPSs.
@@ II System Model
Our study focuses on the downlink (DL) of a cellular network which consists of $M$ terrestrial MBSs and $N$ MBSs installed on a constellation of LEO satellites, making a total of $L$ MBSs. They provide service to $K$ UEs located in the area of study, which is composed of a rural and urban zone. We refer to the overall network bandwidth as $W$, which is distributed by the mobile network operator across the ground and space-based network. We suppose that the network operates within the S band, around 2 GHz, with ground and satellite MBSs using orthogonal, but dynamically adjustable, portions of this band. Throughout this paper, we will use $\mathcal{T}$ and $\mathcal{S}$ to represent the set of terrestrial and satellite MBSs respectively. Moreover, $\mathcal{B}=\mathcal{T}\cup\mathcal{S}=\{1,\cdots,j,\cdots,L\}$ is the complete set of MBSs, and $\mathcal{U}=\{1,\cdots,i,\cdots,K\}$ defines the set of UEs. For the remainder of the paper, we will designate the Hadamard product with $\odot$.
@@ II-A2 Satellite link channel model
It should be noted that components $SF_{ij}$ and $CL$ are function of -and greatly vary with- the LoS conditions, and that the LoS probability for the satellite links are provided in [32, Table 6.6.1-1]. Additionally, it should be highlighted that the elevation angle of the satellite also impacts the quality of the channel link. The elevation angle is the angle between the horizontal plane (the plane parallel to the surface of the Earth at the UE location) and the line of sight to the satellite. If the elevation angle increases, $CL$, $PL_{ij}^{e}$ and $SF_{ij}$ are greatly reduced. Given the cartesian coordinates of a satellite $s$ $\left(x_{s},y_{s},z_{s}\right)$ and a UE $u$ $\left(x_{u},y_{u},z_{u}\right)$, we can compute the elevation angle $\theta_{u}$ as:
$\theta_{u}=\arcsin\left(\frac{z_{s}-z_{u}}{\sqrt{\left(x_{s}-x_{u}\right)^{2}+\left(y_{s}-y_{u}\right)^{2}+\left(z_{s}-z_{u}\right)^{2}}}\right).$ (3)
@@
In (4), $p_{j}$ denotes the transmit power allocated per resource element (RE) at MBS $j$, $\mathcal{I}_{j}$ indicates the set of MBSs interfering with the serving MBS $j$ and $\sigma^{2}$ represents the noise power per RE. Moreover, assuming that MBS $j$ equally shares its total available bandwidth $W_{j}$ between the $k_{j}$ UEs it is serving, we are able to compute the mean throughput for UE $i$ connected to MBS $j$ as follows:
@@
II-C Energy Consumption Model
While [33] provided one of the most widely used models for 4G MBS energy consumption, it is not well suited for 5G MBSs integrating massive MIMO technology. Consequently, we chose the more recent model proposed in [34] which accounts for massive MIMO and carrier shutdown, and thus fits better to our system design. Note that, in the context of our paper, shutdown and sleep mode are considered equivalent, as both refer to states in which a MBS significantly reduces or halts its operations to conserve energy. More generally, it only takes $3$ seconds to shutdown or wake up a MBS, while the shutdown duration may range from tens of seconds to minutes or even hours [35].
The energy consumption of a MBS can be modelled as the sum of multiple components. We denote as the baseline energy consumption, the energy used by the components that are kept active in a shutdown MBS. Then, we denote as the static component, the energy consumption that occurs regardless of the level of the MBS traffic load. The static energy consumption represents the minimum power required to keep essential systems operational and maintain standby readiness. Finally, the dynamic component refers to the load-dependent energy consumption that fluctuates depending on the MBS traffic load. Typically, the dynamic component increases whenever a MBS increases its transmit power or uses additional transmission resources, e.g., more resource blocks. For a MBS $j$, this model can be formulated as:
@@ II-C Energy Consumption Model
where $P_{0}$ represents the baseline energy consumption, $\psi_{j}$ represents the static component and $p_{j}$ accounts for the dynamic consumption of the MBS. Also, $||\cdot||_{0}$ is a binary-valued function equal to $1$ if the transmit power $p_{j}$ is greater than 0.
Regarding the LEO satellite, the total energy consumption can be expressed as the sum of the inherent energy consumption of the LEO satellite, which accounts for altitude adjustments, GPS navigation, and routing operations, plus the energy consumption of the MBS installed on it, as detailed in (6). We suppose that the satellites in the constellation are solar-powered and well-dimensioned. Thus, they can handle both the power requirements needed for an operational satellite and manage the telecom equipment added as payload, based on the growing adoption of real-world projects such as Starlink or Kuiper. The major notations are summarized in Table I.
TABLE I: List of Notations
Parameter Symbol
Carrier frequency $f_{c}$
Total number of MBSs $L$
Total number of UEs $K$
Carrier bandwidth $W$
Set of all MBSs $\mathcal{B}$
Set of all UEs $\mathcal{U}$
Large-scale channel gain for MBS $j$ and UE $i$ $\beta_{ij}$
Elevation angle of the satellite relative to the UE $\theta_{u}$
Noise power per RE [dBm] $\sigma^{2}$
SINR for MBS $j$ and UE $i$ $\gamma_{ij}$
Mean throughput for UE $i$ perceived from MBS $j$ $R_{ij}$
Perceived throughput for UE $i$ $R_{i}$
Energy consumption for MBS $j$ $Q_{j}$
Transmit power per RE at MBS $j$ $p_{j}$
Maximum transmit power per RE at MBS $j$ ${p_{j}}_{\mathrm{max}}$
Vector with transmit power per RE of every MBS $p\in\mathbb{R}^{L}$
Binary matrix for UE-MBS association $X\in\mathbb{R}^{K\times L}$
Fraction of bandwidth allocated to satellite MBSs $\varepsilon$
Coverage Threshold [dBm] $RSRP_{\text{min}}$
Regularization parameter $\lambda$
Hadamard product $\odot$